The ends of latusrectum of parabola are (4, 8) and (4, –8) then equation of parabola is

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
ends of Latus rectum = (4,8) and 4,-8
Therefore a = 4
Equation of parabola = y^2 = 4ax = 4×4 x
y^2= 16x 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)